# RobustSpring: Benchmarking Robustness to Image Corruptions for Optical Flow, Scene Flow and Stereo

### v1

@@ 3.2 Robustness Evaluation Metric
Robustness Metric Concepts. For dense matching, robustness to corruptions is undefined. Metrics exist for adversarial robustness, using the distance between corrupt prediction and either (i) ground-truth [53, 4] or (ii) clean prediction [60, 61, 59]. The latter is preferred for two reasons: First, (i)’s ground-truth comparisons mix accuracy and robustness, which are competing model qualities [60, 71, 68] that should be separate. This competition is intuitive: A model that always outputs the same value is as robust as inaccurate. Likewise, an accurate model varies for any input change and thus is not robust. Second, (ii) separates robustness from accuracy and builds on an established mathematical concept for system robustness [20, 49]: the Lipschitz constant $L^{c}$. It defines robust models as those whose prediction $f$ is similar on clean and corrupt image $I$ and $I^{c}$, relative to their difference. For dense matching, it reads
$L^{c}=\frac{\|f(I)-f(I^{c})\|}{\|I-I^{c}\|}.$ (1)
This robustness formulation is preferable for real-world applications that demand stable scene estimations despite corruptions like snow.
Corruption Robustness Metric. Based on Eq. 1, we quantify model robustness to corruptions. Because RobustSpring’s corrupt images $I_{c}$ deviate from their clean counterparts $I$ by a similar amount, cf. SSIM equalization in Sec. 3.1, we omit the denominator in Eq. 1 and define corruption robustness as distance between clean $f(I)$ and corrupted $f(I^{c})$ predictions with distance metric M:
@@
For similarity to Spring’s evaluation, we use corruption robustness with various metrics M, reporting $R^{c}_{\text{EPE}}$, $R^{c}_{\text{1px}}$ and $R^{c}_{\text{Fl}}$ for optical and scene flow, and $R^{c}_{\text{1px}}$, $R^{c}_{\text{Abs}}$ and $R^{c}_{\text{D1}}$ for stereo. Interestingly, our EPE-based corruption robustness
$R^{c}_{\text{EPE}}=\text{EPE}[f(I),f(I^{c})]=\frac{1}{|\Omega|}\!\sum_{i\in\Omega}\|f_{i}(I)\!-\!f_{i}(I^{c})\|,$ (3)
@@
on image domain $\Omega$ is a generalization of optical-flow adversarial robustness [60] to dense matching and corruptions.
@@
Metric Subsampling. For a benchmark, users should upload robustness results to a web server. To this end, we evaluate on a reduced pixel set by refining Spring’s original subsampling strategy, which retains about 1% of the full data. First, we additionally subsample the set of full-resolution Hero-frames, leaving 0.95%, and then apply 20-fold subsampling, ultimately keeping 0.05% of the full data.
GMFlow MS-RAFT+ FlowFormer GMA SPyNet RAFT FlowNet2 PWCNet
$R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$
Color Brightness 0.33 3.31 1.12 0.33 2.88 1.02 0.68 2.82 1.05 0.36 3.22 1.04 2.72 14.67 8.91 0.92 3.49 1.61 0.45 3.16 1.05 1.04 7.38 3.00
Contrast 0.46 6.71 1.71 0.87 6.69 3.24 0.93 5.48 1.96 0.68 6.43 2.20 8.23 38.90 27.23 1.32 5.73 2.64 1.87 9.26 4.74 2.98 30.07 7.42
Saturate 0.34 3.30 0.96 0.34 2.87 1.03 0.42 2.39 0.88 0.43 3.47 1.18 3.36 17.34 11.31 0.93 3.33 1.47 0.51 3.40 1.10 1.21 9.92 3.68
Blur Defocus 0.53 6.17 1.45 0.51 4.01 1.47 0.55 3.85 1.19 0.56 5.02 2.01 0.57 10.16 1.36 1.03 4.70 2.07 0.53 3.35 1.06 0.98 6.51 2.78
Gaussian 0.66 7.77 1.88 0.58 4.45 1.63 0.63 4.32 1.37 0.62 5.48 2.22 0.76 15.44 2.12 1.10 5.12 2.26 0.60 4.05 1.27 1.11 7.72 3.09
Glass 0.85 20.87 1.82 0.53 4.45 1.37 0.64 4.04 1.17 0.61 5.60 1.91 0.75 16.94 1.36 1.05 5.13 1.97 0.50 3.12 0.96 0.91 5.96 2.47
Motion 1.34 18.35 7.51 1.31 14.06 6.16 1.35 14.03 5.77 1.19 14.40 6.18 2.32 19.55 10.05 2.06 14.33 6.35 1.60 14.07 6.47 1.95 16.25 7.47
Zoom 1.88 35.80 9.90 1.81 21.84 7.13 1.66 22.72 6.77 1.54 23.17 7.16 4.82 46.67 28.37 3.14 22.80 7.61 2.36 24.63 9.04 3.52 50.33 15.64
Noise Gaussian 4.70 57.95 21.67 5.70 35.74 22.12 6.56 27.83 18.30 2.81 24.70 12.96 2.22 42.23 14.88 7.43 27.92 18.99 1.33 11.24 5.06 2.79 26.87 9.89
Impulse 6.64 66.14 28.70 7.39 45.72 29.05 7.33 23.58 14.47 4.08 31.31 18.13 2.92 53.45 20.41 6.51 29.65 18.32 2.37 15.70 7.48 3.57 35.67 14.45
Speckle 3.90 62.01 20.64 4.22 34.96 17.18 5.47 25.52 15.60 5.32 25.22 12.66 1.95 46.32 12.89 6.62 26.05 16.48 1.32 12.57 4.19 2.74 26.83 8.00
Shot 3.52 56.71 17.77 4.36 31.67 17.77 5.75 26.02 16.01 3.15 23.11 11.59 1.86 40.44 11.98 6.74 25.64 17.08 1.16 9.87 3.92 2.59 23.75 7.88
Quality Pixelate 1.96 68.09 18.71 1.60 45.83 6.78 1.48 31.68 2.59 1.11 25.86 1.78 1.22 50.63 2.90 1.65 21.47 2.00 0.77 7.74 0.88 0.92 8.67 2.22
JPEG 3.32 83.54 27.92 2.09 41.69 12.82 2.89 42.62 14.96 1.92 38.70 11.51 2.95 53.97 18.08 3.19 37.72 13.67 2.56 31.00 11.85 2.88 49.15 15.91
Elastic 1.37 40.00 6.89 1.16 32.49 5.54 2.62 35.78 11.01 1.24 27.24 6.40 1.08 34.62 4.77 1.33 19.43 4.78 0.79 16.27 2.12 1.42 28.18 5.47
Weather Fog 0.80 14.42 5.32 0.91 10.32 6.33 0.86 9.66 5.67 0.84 11.21 6.42 5.20 28.15 19.97 1.97 12.01 7.11 1.74 11.77 7.82 16.84 20.96 12.89
Frost 8.20 63.96 29.96 7.38 29.96 21.25 8.18 34.19 23.87 8.13 34.30 22.31 6.97 45.13 30.13 8.37 32.75 21.76 7.22 33.69 21.15 8.27 50.31 27.44
Rain 8.60 64.20 32.72 19.99 36.74 31.22 11.13 33.50 20.83 33.00 43.98 36.18 18.20 68.87 56.38 42.41 38.89 31.99 63.71 48.25 41.15 40.18 73.51 57.05
Snow 3.60 70.60 29.90 4.69 33.21 30.91 7.92 40.20 33.82 5.30 40.82 33.35 12.08 74.27 66.65 7.16 37.04 31.37 39.79 68.67 61.60 39.73 90.80 81.91
Spatter 6.58 67.90 27.09 6.63 28.22 20.24 8.41 40.38 26.92 7.75 36.11 21.81 5.71 48.60 33.82 7.98 30.37 19.87 9.13 45.03 28.99 9.33 65.41 40.19
Average 2.98 40.89 14.68 3.62 23.39 12.21 3.77 21.53 11.21 4.03 21.47 10.95 4.29 38.32 19.18 5.64 20.18 11.47 7.01 18.84 11.09 7.25 31.71 16.44
Median 1.92 48.35 13.83 1.71 29.09 6.95 2.14 24.55 8.89 1.39 23.93 6.79 2.82 41.33 13.88 2.60 22.13 7.36 1.47 12.17 4.90 2.77 26.85 7.94
Clean Error 0.94 10.36 2.95 0.64 5.72 2.19 0.72 6.51 2.38 0.91 7.07 3.08 4.16 29.96 12.87 1.48 6.79 3.20 1.04 6.71 2.82 2.29 82.27 4.89
@@
Table 2: Initial RobustSpring results on corruption robustness of optical flow models, using $R^{c}_{\text{EPE}}$, $R^{c}_{\text{1px}}$ and $R^{c}_{\text{Fl}}$ between clean and corrupted flow predictions. Low values indicate robust models. Clean Error compares clean predictions and ground-truth flows, values from [43].
@@
Robustness Ranking. Because we generate 20 different corruption evaluations per dense matching model, we need a summarization strategy to produce one result per model. Per-model results are ranked based on three strategies: Average, Median, and the Schulze voting method [64]. In contrast to averaging across all 20 evaluations, the median reduces the impact of extreme outliers. The Schulze method provides a holistic, pairwise comparison approach that ranks models based on preference aggregation and was used for prior generalization evaluations in the Robust Vision Challenges. We evaluate their differences in Sec. 4.2
@@ 3.3 Dataset and Benchmark Functionality
Below, we summarize RobustSpring’s corruption dataset and describe its benchmark function. Dataset and benchmark will be at spring-benchmark.org. Fig. 3 shows data samples with stereo, optical flow and scene flow estimates.
RobustSpring Dataset. The final RobustSpring dataset entails 20 corrupted versions of Spring, resulting in 40,000 frames, or 20,000 stereo frame pairs. Each corruption evaluation yields 3960 optical flows (990 per camera & direction), 2000 stereo disparities (1000 per camera) and 3960 additional scene flow disparity maps (990 per camera per direction). We publicly release the RobustSpring test set licensed with CC BY 4.0, but no corrupt training data to discourage corruption finetuning for a fair benchmark.
RobustSpring Benchmark. RobustSpring enables uploading robustness results to a benchmark website for display in a public ranking. To emphasize that robustness and accuracy are two axes of model performance with equal importance [71], we couple RobustSpring with Spring’s established accuracy benchmark. Thus, researchers can report model robustness and accuracy on the same dataset. To maintain Spring’s upload policy, 3 per 30 days, one per hour, each submission receives one robustness upload.

### v2

@@ 3.2 Robustness Evaluation Metric
Definition of Optical Flow. Throughout this work, and consistent with the Spring benchmark (Mehl et al., 2023b), we define optical flow as the true 3D motion of visible surfaces projected into the 2D image plane (Horn and Schunck, 1981; Baker et al., 2011). This definition is standard in classical benchmarks and provides the basis for accuracy evaluation. Alternative formulations, such as apparent motion, exist in the literature, but RobustSpring’s robustness evaluation is agnostic to this choice, since robustness is measured independently of ground truth. A more detailed discussion on the implications of this definition, including the relation between perfect accuracy and perfect robustness, is provided in App. A.1.
Robustness Metric Concepts. For dense matching, robustness to corruptions lacks a standardized evaluation metric. Metrics exist for adversarial robustness, using the distance between corrupt prediction and either (i) ground-truth (Ranjan et al., 2019; Agnihotri et al., 2024c) or (ii) clean prediction (Schmalfuss et al., 2022b; 2023; a). The latter is preferred for two reasons: First, (i)’s ground-truth comparisons mix accuracy and robustness, which are competing model qualities (Schmalfuss et al., 2022b; Tsipras et al., 2019; Taori et al., 2020) that should be separate. This competition is intuitive: A model that always outputs the same value is as robust as inaccurate. Likewise, an accurate model varies for any input change and thus is not robust. Second, (ii) separates robustness from accuracy and builds on an established mathematical concept for system robustness (Hein and Andriushchenko, 2017; Pauli et al., 2022): the Lipschitz constant $L^{c}$. It defines robust models as those whose prediction $f$ is similar on clean and corrupt image $I$ and $I^{c}$, relative to their difference. For dense matching, it reads
$L^{c}=\frac{\|f(I)-f(I^{c})\|}{\|I-I^{c}\|},$ (1)
where the term $\|I-I^{c}\|$ refers to the per-pixel intensity difference between the clean and corrupted images. This robustness formulation is preferable for real-world applications that demand stable scene estimations despite corruptions like snow, and remains valid independent of whether optical flow is defined as true motion or apparent motion. We emphasize that RobustSpring explicitly measures robustness in terms of stability: models are considered robust if their predictions remain consistent under corrupted inputs. Lower robustness scores correspond to higher stability, not improved accuracy. Other definitions of robustness, such as ground-truth-based robustness, are possible and may be more appropriate for different use cases. Our benchmark deliberately focuses on stability-based robustness, providing a complementary axis of evaluation alongside accuracy.
Corruption Robustness Metric. Based on Eq. 1, we quantify model robustness to corruptions. Because RobustSpring’s corrupt images $I_{c}$ deviate from their clean counterparts $I$ by a similar amount, cf. SSIM equalization in Sec. 3.1, we omit the denominator in Eq. 1 and define corruption robustness $R^{c}$ as distance between clean $f(I)$ and corrupted $f(I^{c})$ predictions with distance metric M:
@@
For similarity to Spring’s evaluation, we use corruption robustness with various metrics M, reporting $R^{c}_{\text{EPE}}$, $R^{c}_{\text{1px}}$ and $R^{c}_{\text{Fl}}$ for optical and scene flow, and $R^{c}_{\text{1px}}$, $R^{c}_{\text{Abs}}$ and $R^{c}_{\text{D1}}$ for stereo. Here, EPE denotes the average end-point error, 1px the percentage of pixels with an error exceeding 1 px, Fl the KITTI optical-flow outlier rate, Abs the mean absolute disparity error, and D1 the KITTI disparity outlier rate, see Mehl et al. (2023b) and Menze and Geiger (2015) for more details. Interestingly, our EPE-based corruption robustness
$R^{c}_{\text{EPE}}=\text{EPE}[f(I),f(I^{c})]=\frac{1}{|\Omega|}\!\sum_{i\in\Omega}\|f_{i}(I)\!-\!f_{i}(I^{c})\|$ (3)
@@
on image domain $\Omega$ is a generalization of optical-flow adversarial robustness (Schmalfuss et al., 2022b) to dense matching and corruptions.
@@
Metric Subsampling. For a benchmark, users should upload robustness results to a web server. Given the large number of 20 datasets, data reduction is essential to facilitate evaluations and uploads. To this end, we evaluate on a reduced set of pixels by refining the original subsampling strategy from Spring, which retains about 1% of the full data. First, we additionally subsample the set of full-resolution Hero-frames because they are only kept in the original benchmark for visualization purposes, leaving 0.95%, and then apply 20-fold subsampling, ultimately keeping 0.05% of the full data.
@@
Table 1: Initial RobustSpring results on corruption robustness of optical flow models, using $R^{c}_{\text{EPE}}$, $R^{c}_{\text{1px}}$ and $R^{c}_{\text{Fl}}$ between clean and corrupted flow predictions. Low values indicate robust models. $\varepsilon_{\text{clean}}$ compares clean predictions with ground-truth flow, values from Mehl et al. (2023b).
SEA-RAFT GMFlow MS-RAFT+ FlowFormer GMA SPyNet RAFT FlowNet2 PWCNet
$R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$ $R^{c}_{\text{EPE}}$ $R^{c}_{\text{1px}}$ $R^{c}_{\text{Fl}}$
Color Brightness 0.21 1.65 0.40 0.33 3.31 1.12 0.33 2.88 1.02 0.68 2.82 1.05 0.36 3.22 1.04 2.72 14.67 8.91 0.92 3.49 1.61 0.45 3.16 1.05 1.04 7.38 3.00
Contrast 0.75 3.75 1.51 0.46 6.71 1.71 0.87 6.69 3.24 0.93 5.48 1.96 0.68 6.43 2.20 8.23 38.90 27.23 1.32 5.73 2.64 1.87 9.26 4.74 2.98 30.07 7.42
Saturate 0.16 1.29 0.36 0.34 3.30 0.96 0.34 2.87 1.03 0.42 2.39 0.88 0.43 3.47 1.18 3.36 17.34 11.31 0.93 3.33 1.47 0.51 3.40 1.10 1.21 9.92 3.68
Blur Defocus 0.20 1.47 0.42 0.53 6.17 1.45 0.51 4.01 1.47 0.55 3.85 1.19 0.56 5.02 2.01 0.57 10.16 1.36 1.03 4.70 2.07 0.53 3.35 1.06 0.98 6.51 2.78
Gaussian 0.23 1.79 0.51 0.66 7.77 1.88 0.58 4.45 1.63 0.63 4.32 1.37 0.62 5.48 2.22 0.76 15.44 2.12 1.10 5.12 2.26 0.60 4.05 1.27 1.11 7.72 3.09
Glass 0.22 1.46 0.41 0.85 20.87 1.82 0.53 4.45 1.37 0.64 4.04 1.17 0.61 5.60 1.91 0.75 16.94 1.36 1.05 5.13 1.97 0.50 3.12 0.96 0.91 5.96 2.47
Motion 1.11 13.78 5.82 1.34 18.35 7.51 1.31 14.06 6.16 1.35 14.03 5.77 1.19 14.40 6.18 2.32 19.55 10.05 2.06 14.33 6.35 1.60 14.07 6.47 1.95 16.25 7.47
Zoom 1.83 24.51 8.41 1.88 35.80 9.90 1.81 21.84 7.13 1.66 22.72 6.77 1.54 23.17 7.16 4.82 46.67 28.37 3.14 22.80 7.61 2.36 24.63 9.04 3.52 50.33 15.64
Noise Gaussian 1.49 12.30 4.34 4.70 57.95 21.67 5.70 35.74 22.12 6.56 27.83 18.30 2.81 24.70 12.96 2.22 42.23 14.88 7.43 27.92 18.99 1.33 11.24 5.06 2.79 26.87 9.89
Impulse 2.58 19.01 7.51 6.64 66.14 28.70 7.39 45.72 29.05 7.33 23.58 14.47 4.08 31.31 18.13 2.92 53.45 20.41 6.51 29.65 18.32 2.37 15.70 7.48 3.57 35.67 14.45
Speckle 1.29 12.15 3.61 3.90 62.01 20.64 4.22 34.96 17.18 5.47 25.52 15.60 5.32 25.22 12.66 1.95 46.32 12.89 6.62 26.05 16.48 1.32 12.57 4.19 2.74 26.83 8.00
Shot 1.20 10.69 3.41 3.52 56.71 17.77 4.36 31.67 17.77 5.75 26.02 16.01 3.15 23.11 11.59 1.86 40.44 11.98 6.74 25.64 17.08 1.16 9.87 3.92 2.59 23.75 7.88
Quality Pixelate 0.41 1.88 0.51 1.96 68.09 18.71 1.60 45.83 6.78 1.48 31.68 2.59 1.11 25.86 1.78 1.22 50.63 2.90 1.65 21.47 2.00 0.77 7.74 0.88 0.92 8.67 2.22
JPEG 4.02 34.86 13.57 3.32 83.54 27.92 2.09 41.69 12.82 2.89 42.62 14.96 1.92 38.70 11.51 2.95 53.97 18.08 3.19 37.72 13.67 2.56 31.00 11.85 2.88 49.15 15.91
Elastic 0.56 11.14 1.42 1.37 40.00 6.89 1.16 32.49 5.54 2.62 35.78 11.01 1.24 27.24 6.40 1.08 34.62 4.77 1.33 19.43 4.78 0.79 16.27 2.12 1.42 28.18 5.47
Weather Fog 1.20 11.20 7.41 0.80 14.42 5.32 0.91 10.32 6.33 0.86 9.66 5.67 0.84 11.21 6.42 5.20 28.15 19.97 1.97 12.01 7.11 1.74 11.77 7.82 16.84 20.96 12.89
Frost 7.60 36.86 21.86 8.20 63.96 29.96 7.38 29.96 21.25 8.18 34.19 23.87 8.13 34.30 22.31 6.97 45.13 30.13 8.37 32.75 21.76 7.22 33.69 21.15 8.27 50.31 27.44
Rain 16.73 36.87 25.09 8.60 64.20 32.72 19.99 36.74 31.22 11.13 33.50 20.83 33.00 43.98 36.18 18.20 68.87 56.38 42.41 38.89 31.99 63.71 48.25 41.15 40.18 73.51 57.05
Snow 8.54 60.52 43.81 3.60 70.60 29.90 4.69 33.21 30.91 7.92 40.20 33.82 5.30 40.82 33.35 12.08 74.27 66.65 7.16 37.04 31.37 39.79 68.67 61.60 39.73 90.80 81.91
Spatter 8.93 53.31 30.52 6.58 67.90 27.09 6.63 28.22 20.24 8.41 40.38 26.92 7.75 36.11 21.81 5.71 48.60 33.82 7.98 30.37 19.87 9.13 45.03 28.99 9.33 65.41 40.19
Average 2.96 17.52 9.05 2.98 40.89 14.68 3.62 23.39 12.21 3.77 21.53 11.21 4.03 21.47 10.95 4.29 38.32 19.18 5.64 20.18 11.47 7.01 18.84 11.09 7.25 31.71 16.44
Std. Dev. 4.29 17.98 12.08 2.70 27.91 11.91 4.58 15.54 10.62 3.44 14.37 9.94 7.23 13.67 10.55 4.38 18.35 17.60 9.10 12.55 9.98 15.94 17.93 15.87 11.83 24.43 20.79
Median 1.20 11.68 3.98 1.92 48.35 13.83 1.71 29.09 6.95 2.14 24.55 8.89 1.39 23.93 6.79 2.82 41.33 13.88 2.60 22.13 7.36 1.47 12.17 4.90 2.77 26.85 7.94
$\varepsilon_{\text{clean}}$ (Clean Error) 0.36 3.69 1.35 0.94 10.36 2.95 0.64 5.72 2.19 0.72 6.51 2.38 0.91 7.07 3.08 4.16 29.96 12.87 1.48 6.79 3.20 1.04 6.71 2.82 2.29 82.27 4.89
@@
Robustness Ranking. Because we generate 20 different corruption evaluations per dense matching model, we need a summarization strategy to produce one result per model. Per-model results are ranked based on three strategies: Average, Median, and the Schulze voting method (Schulze, 2018). In contrast to averaging across all 20 evaluations, the median reduces the impact of extreme outliers. The Schulze method provides a holistic, pairwise comparison approach that ranks models based on preference aggregation and was used for prior generalization evaluations in the Robust Vision Challenges. We evaluate their differences in Sec. 4.2.
@@ 3.3 Dataset and Benchmark Functionality
Below, we summarize RobustSpring’s corruption dataset and describe its benchmark function. Fig. 3 shows data samples with stereo, optical flow, and scene flow estimates.
RobustSpring Dataset. The final RobustSpring dataset entails 20 corrupted versions of Spring, resulting in 40,000 frames, or 20,000 stereo frame pairs. Each corruption evaluation yields 3960 optical flows (990 per camera & direction), 2000 stereo disparities (1000 per camera), and 3960 additional scene flow disparity maps (990 per camera per direction). To discourage corruption finetuning and ensure a fair benchmark, we do not provide corrupted training data, only corrupted test data. We separately provide the raw data and a curated dataset for predicting dense matches.
RobustSpring Benchmark. RobustSpring enables uploading robustness results to a benchmark website for display in a public ranking. To emphasize that robustness and accuracy are two axes of model performance with equal importance (Tsipras et al., 2019), we couple RobustSpring with Spring’s established accuracy benchmark. Thus, researchers can report model robustness and accuracy on the same dataset. We provide mock-ups of this integration in App. A.8.